D14610:

Which shows 552 − 452 being evaluated using the difference of squares method? 552 − 452 = (3025 + 2025)(3025 − 2025) = 5,050,000 552 − 452 = 3025 − 2025 = 1,000 552 − 452 = (55 + 45)(55 − 45) = (100)(10) = 1,000 552 − 452 = (55 − 45)2 = 102 = 100

D14610:

i personally think it's c

Phantomdex:

It's C.

Phantomdex:

But we should eliminate the other three options first.

Midnight97:

Yeah it's c bro

KatelynGrace:

yeah c

Phantomdex:

Can you tell me why it wouldn't be the other three?

toga:

The correct answer that shows 552 − 452 being evaluated using the difference of squares method is: 552 − 452 = (55 + 45)(55 − 45) = (100)(10) = 1,000 This is because 552 and 452 are both perfect squares, specifically 552 is equal to (25)2 and 452 is equal to (20)2. Therefore, we can use the formula a2 − b2 = (a + b)(a − b) to rewrite the expression as shown above. Simplifying the expression further, what do we get:

D14610:

d

@phantomdex wrote:
Can you tell me why it wouldn't be the other three?
Well the difference of squares method is a technique used in algebra that manipulates : a difference of squares equation, a² - 6², into the product of the sum and difference of a and b, which is (a + b)(a - b). So, for the expression 552-452, we can apply the difference of squares method to get (55 +45) (55 - 45), which simplifies to (100) (10) = 1000. This corresponds to the third option: 552 45² (55+45) (55 - 45) = (100) (10) = 1,000. The first option is incorrect because it incorrectly calculates the squares of 55 and 45 and incorrectly applies the difference of squares method. The second option is incorrect because it only calculates the squares of 55 and 45 and subtracts them, without applying the difference of squares method. The fourth option is incorrect because it incorrectly applies the difference of squares method resulting in a square of the difference not the difference of squares.